mr.reyes is 20 years older than his son. how old are the father and the son if mr.reyes is 6 times as old as his son?

Answer by ankor@dixie-net.com(22740)

mr.reyes is 20 years older than his son.
r = s+20
:
how old are the father and the son if mr.reyes is 6 times as old as his son?
r = 6s
:
Replace f in the 1st equation with 6s
6s = s + 20
6s - s = 20
5s = 20
s = 20/5
s = 4 yrs is the son
then, obviously:
r = 24 yrs is Reyes age
:
:
See if that works in the 2nd statement
24 = 6(4)